OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5 Numerical Answers Expected! I NEED HELP PLEASSE ???

OpenStudy (callisto):

x − y = −3 -(1) x + 3y = 5 -(2) (1)x3+(2) 3(x − y) + (x + 3y) = 3(−3) + 5 Can you solve it?

OpenStudy (anonymous):

yeah ill type my answer ok and u tell my if is right ... thank

OpenStudy (anonymous):

is -1 .

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

thank u i didnt know what to do first.

OpenStudy (callisto):

Welcome!!~ Do you know how to do it now?

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

can i ask u somethign else

OpenStudy (callisto):

Yes?

OpenStudy (anonymous):

The zip code of a location consists of five digits chosen from the set Z shown below. Z = {0, 1, 2, 3, 4, 5, 6, 7, 8, 9} The set R represents the digits in the zip code for Rubonia. R = {3, 4, 2, 2, 1} The set S represents the digits in one of the zip codes for Sarasota. S = {3, 4, 2, 4, 3} How many prime numbers are in the set ? ~(rUs)

OpenStudy (anonymous):

is 0

OpenStudy (callisto):

You should close this question and start a new post for a new question..

OpenStudy (anonymous):

ok